# Bridging AI and Science: Implications from a Large-Scale Literature Analysis of AI4Science

### v1

@@ 4.2. Uneven Distribution of Bipartite Links
Quantitatively, heavy-tailed degree distributions are observed as visualized in Fig. 2(b-c). Specifically, the degrees for AI method nodes follow a log-normal distribution, while the degree distribution for scientific problem nodes exhibits an even heavier tail, though not as extreme as a power-law distribution. These findings support the idea that a small number of “hubs” are linked to a large variety of AI methods or scientific problems, while many other nodes remain peripheral to these interdisciplinary connections. Heavy-tailed distributions, particularly log-normals, are also commonly observed in other bipartite networks (Vasques Filho and O’Neale, 2018), such as document-term graphs (Wang et al., 2014), and in collaborative networks (Perc, 2010; Sun et al., 2014). This suggests that the structure of the AI4Science bipartite graph reflects a general pattern of interdisciplinary and collaborative work.
Table 12 in the Appendix lists the “hubs” with the highest node degrees, representing the most interconnected scientific problems and AI methods. From the table, we observe that the scientific problems linked to the widest variety of AI methods are primarily computational or data-processing problems, such as: Neural Network Challenges, Statistical Inference, Language and Auditory Processing. These are followed by more domain-specific problems like: Social Media Dynamics, Decision-Making Processes, Market Economics. On the AI methods side, the techniques applied to the greatest number of scientific challenges include general methods such as: Machine Learning, Deep learning, Neural Networks, Classification Methods, Data Analysis, Bayesian Inference. In addition, there are specialized techniques tailored to scientific domains, including: Genomic Analysis, Computational Biology, Protein Design, Molecular Modeling, Material Informatics.
We further break down the node degree distribution based on cluster partitions. Specifically, we calculate the average node degree for well-investigated and under-investigated clusters, as identified earlier in Sec. 4.1. The statistics are listed in Table 2. We observe a higher average degree for well-investigated scientific problems and AI methods. This suggests that under-investigated regions not only have fewer AI4Science publications, but also exhibit a lesser variety, or limited connectivity in the bipartite graph.
@@ 4.3. Discrepancy Between the AI and Science Communities
Table 13 lists the top clusters emphasized by each community. In terms of facilitating scientific discovery with AI, the science community places greater emphasis on challenges such as: Protein Structure and Design, Materials Design, Artificial Intelligence Ethics, and Single-Cell RNA Sequencing; In contrast, the AI community tends to focus more on areas like: Social Media Dynamics, Game Theory, Market Economics, etc. When it comes to applying AI methods to scientific problems, the two communities also show different preferences. In addition to general approaches like Machine Learning and Deep Learning, the science community frequently employs AI techniques specifically tailored for scientific challenges, such as: Genomic Analysis, Protein Design, and Computational Biology. In comparison, the AI community utilizes a broader range of general AI methodologies, covering diverse areas such as: Data Analysis, Game Theory, Causal Inference, Probabilistic Modeling, Reinforcement Learning, Optimization Methods, Generative Models, etc.
@@
The distribution of AI4Science work published in scientific journals and AI conferences in the semantic maps (Fig. 6 in the Appendix) also illustrates the different focuses of these two communities.
@@ 4.4. Findings and Implications
    Different AI and science subdomains exhibit varying degrees of engagement in AI4Science research, leaving a substantial number of scientific problems and AI methods under-investigated in the collaboration context (Sec. 4.1);
@@
    More attention should be directed toward exploring the under-investigated areas, incorporating a broader range of scientific challenges and AI techniques into the AI4Science landscape.
@@
    Efforts should be made to discover new connections between AI and science, uncovering innovative ways to apply AI methods to scientific research.
5. Bridging AI and Science: Through the Lens of Link Prediction
Building on the findings and implications above, we delve deeper into the potentials and challenges of advancing the connections between AI and science from the perspective of link prediction.
@@ 5. Bridging AI and Science: Through the Lens of Link Prediction
Sci $\to$ AI Katz index 0.430 0.335 0.320 0.272 0.093 0.223 0.311 0.450 0.153 0.267 0.315 0.339
Node2vec 0.988 0.836 0.719 0.530 0.259 0.505 0.625 0.765 0.411 0.630 0.669 0.626
LLM+RAG 0.326 0.216 0.200 0.140 0.058 0.096 0.129 0.150 0.099 0.132 0.157 0.145
LLM+Graph 0.541 0.594 0.530 0.437 0.137 0.311 0.419 0.558 0.218 0.408 0.468 0.490
AI $\to$ Sci Katz index 0.485 0.402 0.362 0.291 0.095 0.205 0.262 0.344 0.160 0.271 0.304 0.316
Node2vec 0.997 0.904 0.837 0.717 0.197 0.371 0.470 0.622 0.329 0.526 0.602 0.666
LLM+RAG 0.368 0.293 0.267 0.210 0.036 0.057 0.077 0.098 0.066 0.096 0.120 0.133
LLM+Graph 0.458 0.316 0.300 0.201 0.068 0.113 0.163 0.152 0.118 0.167 0.211 0.173
Table 3. Link prediction results of different models.
Method Link prediction setting Precision Recall F1
@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
Node2vec Sci $\to$ AI Well-investigated 1.000 0.980 0.922 0.799 0.115 0.295 0.417 0.606 0.207 0.453 0.574 0.689
Under-investigated 0.960 0.726 0.535 0.310 0.392 0.734 0.841 0.928 0.557 0.730 0.654 0.465
AI $\to$ Sci Well-investigated 0.996 0.956 0.916 0.807 0.126 0.293 0.414 0.616 0.224 0.448 0.571 0.699
Under-investigated 0.978 0.651 0.481 0.284 0.560 0.836 0.918 0.978 0.712 0.732 0.631 0.440
LLM+RAG Sci $\to$ AI Well-investigated 0.422 0.271 0.247 0.189 0.031 0.054 0.074 0.102 0.057 0.090 0.113 0.132
Under-investigated 0.240 0.178 0.157 0.111 0.081 0.149 0.182 0.224 0.121 0.162 0.169 0.148
AI $\to$ Sci Well-investigated 0.344 0.259 0.267 0.206 0.027 0.053 0.076 0.098 0.050 0.088 0.118 0.133
Under-investigated 0.188 0.115 0.089 0.066 0.090 0.140 0.159 0.167 0.122 0.126 0.114 0.095
@@ 5.1. Methodology
We split the data into training and test sets based on publication dates. The training set includes papers published between 2014 and 2022 (5,829 AI4Science publications), which are used for training models, retrieving data points for generation augmentation, and identifying well/under-investigated areas, as illustrated in Fig. 3. The testing set consists of publications from 2023 and 2024 (1,195 AI4Science publications), serving as the ground truth for evaluating link prediction models. The statistics are provided in Table 14 in the Appendix. It is important to note that the analyses prior to this section involve the full AI4Science dataset, spanning the period from 2014 to 2024.
@@
We employ two types of models to predict links between scientific problems and AI methods. The first category includes conventional bipartite link prediction models (Özer et al., 2024), such as those based on the Katz index (Katz, 1953) and node2vec embeddings (Grover and Leskovec, 2016). Formally, given a source node $P_{i}$ (or $M_{i}$) in the bipartite graph, we predict $K$ potential target nodes for linking. We then compare the predicted set with the ground truth target node set and report Precision, Recall, and F1 scores @K (Al Hasan et al., 2006; Yang et al., 2015). The bipartite graph is constructed and the models are trained using only the training data, with evaluation performed on the testing data. Please note that existing links may also be counted in constructing the ground truth set, because a new AI4Science work may reuse an existing connection between a scientific problem and an AI method.
In addition to conventional link prediction, we utilize a large language model (LLM) for generative predictions. Specifically, given a scientific problem $p_{i}$ from cluster $P_{i}$, we generate potential AI methods $m_{i}$ that could be applied. These generated texts are then mapped to cluster labels $M_{i}$ through embedding similarity matching. Similarly, given an AI method $m_{i}$, we can generate potential scientific problems $p_{i}$ that it could address. We use the OpenAI gpt-3.5-turbo-0125 model as the generator. Since this model’s knowledge is current up to September 2021, there is no risk of data leakage. The prediction process is further enhanced by retrieving similar papers from the training set using retrieval-augmented generation (RAG) (Lewis et al., 2020).
While we can use large language models for generative predictions leveraging embedding similarity matching, the large language model is only exposed to a small neighborhood of literature through RAG and has no access to the higher-level linked bipartite graph structure. Thus, we include results where the entire graph (training links) is provided to the model. Given a scientific problem cluster $P_{i}$, all possible AI method clusters $M$ and past links between the clusters, we directly generate AI method cluster labels $M_{i}$. Similarly, given a AI methods cluster $M_{i}$, all possible scientific problem clusters $P$ and past links between the clusters, we directly generate scientific problem cluster labels $P_{i}$.
For more detailed descriptions of the models and evaluation metrics, please refer to Appendix C.
@@ 5.2. Overall Link Prediction Results
Table 3 presents models’ performances on the test data, where conventional link prediction methods show strong results in both Precision and Recall. Particularly, while both Katz and node2vec rely on suggesting similar nodes to link, node2vec outperforms Katz, suggesting the underlying mechanism connecting scientific problems and AI methods goes beyond homophily in the local neighborhood and potentially considers the broader structures of the AI4Science network.
Notably, the LLM+RAG approach lags behind conventional link prediction methods, emphasizing the challenge of generative link prediction at the individual paper level. A potential explanation is that, unlike node-level link prediction in a bipartite graph, LLMs struggle to obtain a comprehensive global view of the AI4Science landscape, even with the aid of RAG. This implies the benefit of considering the global bipartite graph structures and the opportunity to build knowledge graphs in the AI4Science domain to help LLMs better understand the broader context.
This is further elaborated by the results for LLM+Graph, where the model is provided with global bipartite graph structure. LLM+Graph outperforms LLM+RAG, suggesting that LLM benefits from bipartite linked graph information. Moreover, LLM+Graph achieves comparable results with Katz index, which implies LLM’s ability to understand graph-level features and make informed link predictions.
Nevertheless, LLM+RAG offers a unique advantage: it can generate detailed ideas in texts for AI4Science applications. Table 19 in the appendix reports evaluation results on text generation metrics, where LLM+RAG demonstrates notable results.
5.3. Prediction in the Well- and Under-investigated Regions
Fig. 3 reveals the presence of well- and under-investigated areas in both the scientific problem and AI method spaces. Echoing Implication (I1), we comparatively explore the potentials and challenges in facilitating interdisciplinary work for these regions.
Formally, we refer to scientific problem clusters or AI method clusters above the regression lines as well-investigated areas, while those that fall below the lines are the under-investigated regions. Clusters that reside within the 95% confidence intervals of the regression results are excluded from this analysis.
Table 4 compares the link prediction results for well-investigated and under-investigated clusters. In both categories of link prediction methods, we observe a similar trend: Precision scores are generally higher for well-investigated scientific problems and AI methods, even as K increases. In contrast, Recall is relatively higher for under-investigated clusters. These results are consistent with the node degree statistics in Table 2, where the average degree of under-investigated nodes is much lower, making it easier for the models to predict links, resulting in higher Recall. For instance, node2vec can retrieve nearly all ground truth links with K=10.
Additionally, in the under-investigated partition, the rapid decline in Precision as K increases suggests that link prediction models are capable of discovering novel links that have not been identified by human researchers in the test data. This highlights great opportunities for expanding the connections between AI methods and scientific challenges.
Reference Method # of novel (unseen) links
@@ 5.3. Prediction in the Well- and Under-investigated Regions
Train Node2vec 254 820 1,369 2,722
LLM+RAG 870 2,502 3,983 6,775
Train+Test Node2vec 4 396 899 2,203
LLM+RAG 760 2,308 3,748 6,505
Table 5. Number of novel links discovered by models compared with the training set and the whole dataset. The test data introduced 554 new links compared to the training data.
@@ 5.4. Discovering Novel Links
Previous discussions highlight the capability of link prediction models to uncover novel links. Building on this, we further investigate the newly discovered connections in response to Implication (I2).
Compared to the training data up to 2022, human researchers identified 554 new links between scientific problems and AI methods during 2023 and 2024. In contrast, link prediction methods have demonstrated the potential to discover even more connections, especially as K increases, as shown in Table 5.
Remarkably, the LLM+RAG approach demonstrates strong potential in generating new links, even with a relatively small number of predictions per data point (e.g., @1). This observation aligns with recent research suggesting that LLMs are capable of proposing novel research ideas (Si et al., 2024). Many of the predicted links involve applying Machine Learning, Deep Learning, and various types of Neural Networks (e.g., GNN, RNN, GAN), as well as Reinforcement Learning, to address a wide range of scientific challenges.
In comparison, conventional link prediction methods like node2vec, which leverage cluster-level bipartite graph structures, exhibit high accuracy in identifying new links in the test data but show lower novelty compared to LLM+RAG. For instance, in the @1 predictions, the node2vec model successfully identifies 250 of the 254 new links (98.4%) that were also discovered by human researchers. The four remaining links, not identified by human researchers, include: Cardiac Pathophysiology – Matrix Factorization, Mycobacterium Tuberculosis – Contrastive Learning, Plant Stress Responses – Topological Data Analysis, and Prostate Cancer – Ensemble Learning. Some of these links have since been developed into publications (Wang et al., 2019; Tse et al., 2020; Palande et al., 2023; Patel et al., 2024), highlighting the model’s ability to suggest valuable connections ahead of human exploration.

### v2

@@ 3.2. Uneven Distribution of Bipartite Links
Quantitatively, heavy-tailed degree distributions are observed as visualized in Fig. 4(b-c). Specifically, the degrees for AI method nodes follow a log-normal-like distribution, while the degree distribution for scientific problem nodes exhibits an even heavier tail, though not as extreme as a power-law distribution. These findings support the idea that a small number of “hubs” are linked to a large variety of AI methods or scientific problems, while many other nodes remain peripheral to these interdisciplinary connections. Heavy-tailed distributions, particularly log-normals, are also commonly observed in other bipartite networks (Vasques Filho and O’Neale, 2018), such as document-term graphs (Wang et al., 2014), and in collaborative networks (Perc, 2010; Sun et al., 2014). This suggests that the structure of the AI4Science bipartite graph reflects a general pattern of interdisciplinary and collaborative work.
Table 13 in Appendix lists the “hubs” with the highest node degrees, representing the most interconnected scientific problems and AI methods. From the table, we observe that the scientific problems linked to the widest variety of AI methods are primarily computational or data-processing problems, such as: ‘Neural Network Challenges’ and ‘Statistical Inference’. These are followed by more domain-specific problems like: ‘Urban Traffic Management’, ‘Electronic Health Data Challenges’, ‘Visual Perception and Neuroscience’, ‘Market Pricing and Allocation’, ‘Single-cell RNA Sequencing’, ‘Soft Robotics’, and ‘Drug Discovery’.
On the AI methods side, the techniques applied to the greatest number of scientific challenges include general methods such as: ‘Machine Learning’, ‘Deep learning Models’, ‘Neural Networks’, ‘Classification Methods’, ‘Data Analysis’. In addition, there are specialized techniques tailored to scientific domains, including: ‘Genomic Analysis Methods’, ‘Computational Drug Screening’, ‘Molecular Dynamics Simulations’, ‘Protein Structure Design’, and ‘Molecular Structural Design’.
Node type Partition Avg. Deg.
Scientific problem clusters (n=390) All nodes 13.0
Well-explored 41.3
Under-explored 5.0
AI method clusters (n=355) All nodes 14.3
Well-explored 19.1
Under-explored 8.1
Table 2. Average degree of nodes in the bipartite graph.
We further break down the node degree distribution based on cluster partitions. Specifically, we calculate the average node degree for well-explored and under-explored clusters, as identified earlier in Sec. 3.1. The statistics are listed in Table 2. We observe a higher average degree for well-explored scientific problems and AI methods. This suggests that under-explored regions not only have fewer AI4Science publications, but also exhibit a lesser variety, or limited connectivity in the bipartite graph.
@@ 3.3. Discrepancy Between the AI and Science Communities
Table 14 in Appendix lists the top clusters emphasized by each community. In terms of facilitating scientific discovery with AI, the science community places greater emphasis on challenges such as: ‘Single-cell RNA Sequencing’, ‘Mechanical Properties of Materials’, ‘Protein Design and Function Prediction’, ‘Cancer Detection’, and ‘Genetic Variants and Traits’; In contrast, the AI community tends to focus more on areas like: ‘Urban Traffic Management’, ‘Electronic Health Data Challenges’, ‘Statistical Inference’, ‘Market Pricing and Allocation’, ‘Soft Robotics’, etc.
When it comes to applying AI methods to scientific problems, the two communities also show different preferences. In addition to general approaches like ‘Machine Learning’ and ‘Deep Learning Models’, the science community frequently employs AI techniques specifically tailored for scientific challenges, such as: ‘Genomic Analysis Methods’, ‘Molecular Dynamics Simulations’, and ‘Computational Drug Screening’. In comparison, the AI community utilizes a broader range of general AI methodologies, covering diverse techniques such as: ‘Reinforcement Learning’, ‘Attention Mechanisms’, ‘Policy Gradient Methods’, ‘Multitask Learning’, ‘Spatio-Temporal Analysis’, ‘Causal Inference’, ‘Self-Supervised Learning’, and ‘Unsupervised Learning’.
@@
The distribution of AI4Science work published in scientific journals and AI conferences in the semantic maps (Fig. 7 in Appendix) also illustrates the different focuses of these two communities.
@@ 3.4. Findings and Implications
    Different AI and science subdomains exhibit varying degrees of engagement in AI4Science research, leaving a substantial number of scientific problems and AI methods underexplored in the collaboration context (Sec. 3.1);
@@
    More attention should be directed toward exploring the under-explored areas, incorporating a broader range of scientific challenges and AI techniques into the AI4Science landscape.
@@
    Efforts should be made to discover new connections between AI and science, uncovering innovative ways to apply AI methodologies to scientific research.
In the following section, we present preliminary explorations of addressing these implications through link prediction.
4. Bridging AI and Science:
Through the Lens of Link Prediction
Building on the findings and implications above, we dive deeper into the potentials and challenges of advancing the connections between AI and science from the perspective of link prediction.
@@ 4. Bridging AI and Science:
Sci $\to$ AI Katz index 0.115 0.108 0.086 0.076 0.015 0.035 0.052 0.104 0.026 0.053 0.065 0.088
Node2vec 0.198 0.185 0.168 0.129 0.033 0.098 0.138 0.197 0.056 0.128 0.151 0.156
LLM (Cluster) 0.352 0.300 0.238 0.159 0.053 0.101 0.134 0.176 0.093 0.151 0.171 0.167
LLM (Paper) 0.282 0.227 0.200 0.156 0.058 0.106 0.129 0.155 0.096 0.145 0.157 0.155
AI $\to$ Sci Katz index 0.094 0.120 0.110 0.102 0.010 0.028 0.041 0.079 0.018 0.045 0.060 0.089
Node2vec 0.096 0.108 0.110 0.109 0.013 0.042 0.058 0.104 0.023 0.060 0.076 0.106
LLM (Cluster) 0.201 0.168 0.143 0.129 0.016 0.042 0.053 0.092 0.030 0.068 0.077 0.107
LLM (Paper) 0.377 0.323 0.307 0.296 0.065 0.104 0.127 0.163 0.111 0.158 0.180 0.211
Table 3. Link prediction results of different models. Bold numbers highlights the highest performances in each scenario.
Link prediction Precision Recall F1
Setting Method Region @1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
Sci $\to$ AI LLM (Paper) Well-explored 0.293 0.259 0.230 0.184 0.025 0.059 0.073 0.105 0.046 0.095 0.111 0.134
Under-explored 0.250 0.195 0.171 0.129 0.111 0.192 0.224 0.247 0.153 0.193 0.194 0.169
LLM (Cluster) Well-explored 0.360 0.411 0.353 0.228 0.030 0.082 0.123 0.164 0.055 0.137 0.182 0.191
Under-explored 0.194 0.125 0.089 0.051 0.068 0.122 0.147 0.182 0.101 0.124 0.111 0.079
AI $\to$ Sci LLM (Paper) Well-explored 0.403 0.340 0.334 0.332 0.045 0.078 0.096 0.126 0.080 0.127 0.149 0.182
Under-explored 0.364 0.301 0.267 0.236 0.151 0.222 0.264 0.311 0.214 0.255 0.265 0.269
LLM (Cluster) Well-explored 0.328 0.251 0.183 0.185 0.020 0.047 0.053 0.102 0.038 0.079 0.082 0.132
Under-explored 0.076 0.057 0.074 0.102 0.025 0.051 0.051 0.037 0.038 0.054 0.060 0.055
@@ 4.1. Methodology
We split the data into training and test sets based on publication dates. The training set includes papers published between 2014 and 2022 (6287 AI4Science publications), which are used for training models, retrieving data points for generation augmentation, and identifying well/under-explored areas, as illustrated in Fig. 3. The testing set consists of publications from 2023 and 2024 (1225 AI4Science publications), serving as the ground truth for evaluating link prediction models. The statistics are provided in Table 15 in Appendix. Note that the analyses prior to this section involve the full AI4Science dataset, spanning the period from 2014 to 2024.
@@
We employ two types of models to predict links between scientific problems and AI methods. The first category directly make predictions on the cluster level with conventional bipartite link prediction models (Özer et al., 2024), such as those based on the Katz index (Katz, 1953) and node2vec embeddings (Grover and Leskovec, 2016) or large language models (LLMs). Formally, given the cluster-level bipartite graph constructed on the training data and a source node $P_{i}$ (or $M_{i}$), the model predicts $K$ target nodes that could be potentially linked to the source node. We then compare the predicted nodes with the ground-truth target node set in the test data and report Precision, Recall, and F1 scores @K (Al Hasan et al., 2006; Yang et al., 2015). Note that existing links may also appear in the ground-truth set, because a new AI4Science work may repeat an existing connection between a scientific problem and an AI method. Specifically, when using LLM for link prediction, the bipartite links to the source nodes are provided in prompts, and the model is referred to as ‘LLM (Cluster)’.
In addition, we also utilize LLMs for paper-level generative predictions. Specifically, given a scientific problem $p_{i}$ from cluster $P_{i}$, we prompt the LLM to generate potential AI methods $m_{i}$ that could be applied. These generated method descriptions are then mapped to cluster labels $M_{i}$ through embedding similarity. Similarly, given an AI method $m_{i}$, we prompt the LLM to generate potential scientific problems $p_{i}$ that it could address. The prediction process is further enhanced by retrieving similar papers from the training set using retrieval-augmented generation (RAG) (Lewis et al., 2020). The model is referred to as ‘LLM (Paper)’.
We use OpenAI’s gpt-4o-2024-08-06 as the LLM model and report its performance in the main text. Additionally, we also tested the gpt-3.5-turbo-0125 model as the generator, given that its knowledge is limited to September 2021, ensuring no risk of data leakage. Our results show that the performance of LLM (Paper) remains consistent, suggesting a low risk of data leakage. However, the performance of LLM (Cluster) declines, likely due to GPT-3.5’s limited ability of reasoning and processing long contexts. For more detailed descriptions of the models and evaluation metrics, please refer to Appendix C.
@@ 4.2. Overall Link Prediction Results
Table 3 presents the models’ performances on the test data, where LLM-based link prediction methods outperform conventional approaches in most settings, highlighting their strong potential for predicting AI4Science research directions.
[…]
For conventional link prediction methods, while both Katz and node2vec rely on suggesting similar nodes to link to, node2vec outperforms Katz, suggesting the underlying mechanism connecting scientific problems and AI methods goes beyond homophily in the local neighborhood and potentially considers the broader structures of the AI4Science network. This finding further highlights the potential utility of the AI4Science bipartite graph.
4.3. Well- and Under-Explored Regions
Fig. 3 reveals the presence of well- and under-explored areas in both the scientific problem and AI method spaces. Echoing Implication (I1), we comparatively explore the potentials and challenges in facilitating interdisciplinary work for these regions.
Formally, we refer to scientific problem clusters or AI method clusters above the regression lines as well-explored areas, while those that fall below the lines are the under-explored regions. Clusters that reside within the 95% confidence intervals of the regression results are excluded from this analysis. Note that for this experiment, we use only the publications in the training split to decide the well- and under-explored regions.
Table 4 compares the link prediction results for well-explored and under-explored clusters. In both categories of link prediction methods, we observe a similar trend: Precision scores are generally higher for well-explored scientific problems and AI methods, even as K increases. In contrast, Recall is relatively higher for under-explored clusters. These results are consistent with the node degree statistics in Table 2, where the average degree of under-explored nodes is much lower, making it easier for the models to predict links, resulting in a higher Recall.
Additionally, in the under-explored partition, the rapid decline in Precision as K increases suggests that link prediction models are capable of discovering novel links that have not been actually explored by researchers in the test period. This highlights great opportunities for using “AI” to facilitate the expansion of the connections between AI methods and scientific challenges.
Method # of novel (unseen) links
@@ 4.3. Well- and Under-Explored Regions
Node2vec 8 132 432 1623
LLM (Cluster) 243 867 1,523 3,209
LLM (Cluster) 430 1,650 2,999 6,963
LLM (Paper) 639 2,291 3,771 8,276
Table 5. Number of novel links discovered by models compared with the training set (data up to 2022). The publications during 2023 and 2024 introduced 683 new links. For a fair comparison, in LLM (Cluster) we control the number of API calls to match that of LLM (Paper).
@@ 4.4. Discovering Novel Links
Inspired by the capability of link prediction models to uncover novel links, we further explore the newly discovered connections in response to Implication (I2).
Compared to the training data up to 2022, publications during 2023 and 2024 actually identified 683 new links between scientific problems and AI methods. In contrast, link prediction methods exhibit the potential to discover even more connections, especially as K increases, as shown in Table 5. Remarkably, the LLM-based approaches exhibit strong potential in generating new links, even with a relatively small number of predictions per data point (e.g., @1). This observation aligns with recent research suggesting that LLMs are capable of proposing novel research ideas (Si et al., 2024). Interestingly, LLM (Paper) demonstrates a stronger capability in discovering novel links compared to LLM (Cluster), even when the number of API calls is controlled. This may be due to the greater diversity of paper-level inputs, leading to more varied outputs.
Many of the predicted links have already materialized into publications, such as applying ‘Graph Neural Networks’ to ‘Protein Design and Function Prediction’ (Meller et al., 2023), using algorithms for ‘Molecular Dynamics Simulations’ to tackle challenges in ‘Polymer Dynamics and Chirality’ (Kuenneth and Ramprasad, 2023) and ‘Earth’s Mantle Dynamics’ (Li et al., 2023), or leveraging ‘Multimodal Learning’ for ‘Visual Perception and Neuroscience’ (Choi et al., 2023). Link prediction also highlights numerous unexplored but promising directions, such as using ‘Reinforcement Learning’ for ‘Carbon Emissions Mitigation’, applying ‘Graph Neural Networks’ to predict ‘RNA-Protein Interactions’, and using ‘Explainable AI’ to address challenges in ‘Artificial Intelligence Ethics‘.
5. Related Work
5.1. Qualitative Literature Reviews of AI4Science
While these reviews provide valuable insights, they often rely on interdisciplinary experts to manually identify and categorize scientific problems and potential AI methods, limiting their scope. Such an approach often could not offer a comprehensive and data-driven overview of the broader AI4Science landscape.
5.2. Quantitative Literature Analysis
In addition to heuristic and qualitative literature reviews, a few recent papers from the Science of Science community have focused on quantitatively analyzing the AI4Science publications. For instance, Gao and Wang (2023) discovered that AI use is prevalent across various scientific disciplines, and papers incorporating AI tend to receive a citation advantage. Duede et al. (2024) explored AI’s engagement in different fields, uncovering an “oil-and-water” phenomenon – while AI-engaged work is spreading across disciplines, it does not integrate well with other work within each specific field.
Although these studies provide valuable insights into AI4Science, they primarily focus on areas where AI is already applied to the scientific problems, overlooking the broader and underexplored parts of the landscape. Moreover, these approaches often rely on pre-defined keywords or n-grams to tag AI4Science research, leading to a taxonomy that provides only a coarse representation of the AI and science space. This method fails to capture the actual semantics and nuances of the language, ultimately limiting the depth of analysis.
By contrast, our study provides a more comprehensive view by analyzing literature from both AI and scientific venues over the past decade, encompassing not only AI4Science but also non-AI4Science publications. We leverage LLMs to extract detailed textual descriptions, enabling a nuanced and scalable data-driven categorization that facilitates deeper analysis. Our depiction of the landscape, quantitative analysis of discrepancies, and link prediction experiments offer fresh insights into the challenges and opportunities in bridging AI and science.